A sum of $2000 is invested at an annual interest rate of 5%, compounded monthly. Find the balance in the account after 6 years. (Round your answer to the nearest cent.)
Found 2 solutions by rfer, addingup:
Answer by rfer(16322)

FV=2000(1+0.05/12)^(12*6)
FV=2000(1.00417)^72
FV=2000(1.34934)
FV=2698.68

Answer by addingup(3677)

2000(1+(0.05/12))^(12*6)
= 2000(1.3490) = 2698